OpenStudy (anonymous):

the lcd of 5/12 and -9/16

OpenStudy (anonymous):

@sloppycanada

OpenStudy (sloppycanada):

What two numbers do they have in common?

OpenStudy (sloppycanada):

Think of above their current base. Seeing as you don't want to simplify.

OpenStudy (anonymous):

is the answer 24

OpenStudy (wolf1728):

The lcd is not 24

OpenStudy (anonymous):

what is it

OpenStudy (wolf1728):

It is 48 So, you will have to convert 5/12 and -9/16 into 48ths.

OpenStudy (wolf1728):

You know how to do that?

OpenStudy (wolf1728):

In order to convert 5/12 into 48ths we multiply by 4 5/12 = 20/48 In order to convert -9/16 into 48ths we multiply by 3 -9/16 = -27/48